A particle starts at x=0 and moves along the x-axis with velocity v(t)=2t+1 for time t is less than or equal to 0. Where is the particle at t=4?
Answer #1
the question makes no sense. It starts at time zero, and there is no definition of velocity after that.
answered by: chuchuchulu
Answer #2
I copied the problem out of my textbook
answered by: ivonna
Answer #3
I think you misstated the problem. Is it really
"A particle starts at x=0 and moves along the x-axis with velocity v(t)= 2t+1 for time t GREATER than or equal to 0. Where is the particle at t=4?" ?

If so, position is the integral of velocity from t=0 to t = 4.
That integral is the change in t^2 + t from t = 0 to 4, which is 20.

You can get the same answer by multiplying the average velocity for the interval, which is 5, but the time interval (4).
answered by: megajolt
Answer #4
WHOA> STOP.

the v(t) is only valid for t Less than t=0. So the answer is unable to determine.
answered by: shawntae
Answer #5
oh okay well Thank You for you help I will ask my teacher how to solve this tomorrow
answered by: shakiela
Answer #6
I assume you are in calculus.

Postition=INTEGRAL v(t) dt=INT (2t+1)dt
= t^2 + t

Put in t=4 and compute.
Answer #7
What is your subject?
Answer #8
oh okay Thank You
answered by: Andrea Williams
Answer #9
not basically 20, it is 20. Units were not specified.
answered by: jeep
Answer #10
The problem asks for postion at time 4, which is outside the time given for the velocity expression, if you typed it correcty. One needs the velocity function for time zero to somewhat beyond 4 to calculate postion.
answered by: Electricity:providing energy
Answer #11
so t less than or equal to 0 is invaild
answered by: Alex Pedraza
Answer #12
barbara, you originally asked for position at time 4 for a function that has a certain velocity function before time zero. One cant calculate postition for time 4 unless there is some indication of what it was doing between time zero and time 4.
answered by: giddy
Answer #13
Calculus
answered by: mamta
Answer #14
the velocity v(t)= 2t + 1
answered by: Jay Jay
Answer #15
so my answer would basically be 20